OpenStudy (anonymous):

Mr. Crosson selects two students from his class to present their reports on Spain. He places the names of each student in a box and picks one at a time. The class has 8 boys and 10 girls. What is the probability that Mr. Crosson will pick two boys in a row if he does not put the first name back in the box? Round your answer to the nearest percent Can someone help me out ?

OpenStudy (anonymous):

i know once you mulitply the fractions you get 56/306 , do i simplify to find the percent ?

OpenStudy (anonymous):

8 over 10

OpenStudy (anonymous):

and yes

OpenStudy (anonymous):

jc?

OpenStudy (anonymous):

yes ?

OpenStudy (kropot72):

The probability of picking a boy in the first draw is 8/18. Having picked a boy in the first draw, the probability of picking a boy in the next draw is 7/17. \[P(2\ boys\ sequentially)=\frac{8}{18}\times\frac{7}{17}=you\ can\ calculate\] It is usual to either simplify the fraction or to express the answer as a decimal number to a specified number of places.

OpenStudy (anonymous):

I got 18.3 percent ?

OpenStudy (kropot72):

In the question it asks for the answer to be expressed to the nearest percent. Therefore the answer is 18%.

OpenStudy (anonymous):

thanks bro

OpenStudy (kropot72):

You're welcome :)

OpenStudy (anonymous):

Margo and Carrie are trying to go to Idaho for Dawn's wedding. The probability of finding a flight for less than $400 is 1/7. The probability of Carrie making enough money babysitting to pay for her flight is 3/5. The probability of the girls being able to take time off from work is 1/2. What is the probability that they will make it to their friend's wedding? Want to help me out with this one, i have online classes so i practically have to teach myself these forumals without 1 mistake or i fail lol.

OpenStudy (kropot72):

This second question is not clearly expressed. However it appears that it is certain that Margo can afford to pay for the flight even if the cost is more than $400. Therefore the probability of both women getting to the wedding is dependent on Carrie's being able to pay for the flight and both women getting time off, which is \[\frac{1}{7}\times\frac{3}{5}\times\frac{1}{2}=you\ can\ calculate\]

OpenStudy (anonymous):

3/70 ?